OpenStudy (anonymous):

prove that a^2+b^2/2 greater than or equal to (a+b/2)^2

OpenStudy (anonymous):

no u cannot do so

OpenStudy (anonymous):

sorry i didn't read carefully

OpenStudy (anonymous):

hint given is (a+!)^2+(b+1)^2 is greater than equal to 0

OpenStudy (anonymous):

Is it \(\large \frac{a^2+b^2}{2}\ge (\frac{a+b}{2})^2\)?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

can u solve it...

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

uh....im probably way off but.... if everything is broken down, 2ab is less than or equal to zero and because the equation turns out like this:\[(a^2+b^2)/2\ge (a^2+2ab+b^2)/4\]

OpenStudy (anonymous):

14yamaka in my book he has used (a+1)^2+(b+!)^2>/=0

OpenStudy (anonymous):

is it an exclamation mark after b+?

OpenStudy (anonymous):

sorry it is 1

OpenStudy (anonymous):

Note that \(\large (\frac{a+b}{2})^2=\frac{a^2}{4}+\frac{b^2}{4}+\frac{ab}{2}\). The inequality can then be written as: \[a^2+b^2\ge \frac{a^2}{2}+\frac{b^2}{2}+ab=a^2+b^2-(\frac{a^2}{2}+\frac{b^2}{2}-ab).\] From here it would enough to show that \(-(\frac{a^2}{2}+\frac{b^2}{2}-ab)\le 0\) or \((\frac{a^2}{2}+\frac{b^2}{2}-ab)\ge 0.\) \[(\frac{a^2}{2}+\frac{b^2}{2}-ab)=\frac{1}{2}(a-b)^2\ge 0.\]

OpenStudy (anonymous):

Ask if you didn't get any of the steps.

OpenStudy (anonymous):

explain the second step

OpenStudy (anonymous):

I multiplied both sides by 2 first then I wrote \[\frac{a^2}{2}+\frac{b^2}{2}+ab=a^2-\frac{a^2}{2}+b^2-\frac{b^2}{2}+ab=a^2+b^2-(\frac{a^2}{2}+\frac{b^2}{2}-ab).\] You can then write the inequality as \[a^2+b^2\ge a^2+b^2-(\frac{a^2}{2}+\frac{b^2}{2}-ab) \implies -(\frac{a^2}{2}+\frac{b^2}{2}-ab)\le 0 \implies \cdots \]

OpenStudy (anonymous):

thanks for helping anwar